# Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap

### v1

@@ Abstract
A participant wearing Apple Vision Pro performs spatial decision making tasks using two different visualizations: AR See-Through (Left) and 2D Minimap (Right). Both visualizations present AI-suggested targets, but differ in representations. The AR See-Through approach embeds information directly into the participant’s field of view, while the 2D Minimap offers a top-down abstraction.
Introduction
In high-stakes, time-critical scenarios – such as emergency evacuation [89, 50], first-responder prioritization [96], and crisis management [40] – decision-makers must rapidly select among multiple spatial targets (e.g., exits, individuals requiring assistance, or areas to secure). We refer to these decision-making tasks, which require integrating information about spatial environments with external data, as Spatial Decision-Making. Making the wrong choice in these tasks can waste valuable resources or endanger lives [31]. Recent advancements in indoor sensing and artificial intelligence (AI) offer new avenues for supporting such decisions by visualizing real-time situational data and AI suggestions in spatial context [59, 44].
However, the way this information is presented can significantly impact user behavior and decision quality [24]. Existing methods [45, 92] commonly employ 2D maps to display data and AI suggestions, requiring users to mentally map and connect these digital cues onto physical spaces. According to spatial cognition research [38], this process – known as reference-frame translation – can impose significant cognitive load. Under time pressure, users may struggle to critically assess AI outputs, resulting in inappropriate reliance [74]: either trusting flawed AI suggestions or rejecting correct ones. This concern is especially pressing in emergency contexts, where every second counts.
Augmented Reality (AR) promises a more intuitive approach by directly visualizing situational data and AI suggestions onto the physical environment. These visualizations – often referred to as embedded visualizations [87, 42, 8, 69] – can reduce cognitive load by eliminating the need for constant reference-frame translation. Intuitively, embedded visualizations might foster more deliberate and appropriate reliance on AI. Yet whether embedded visualizations genuinely mitigates inappropriate reliance on AI in spatial decision-making remains unclear, and little existing work has examined how embedded visualizations in AR affects human-AI reliance in spatial tasks. As AR and AI continue to mature and increasingly integrate [32], this work aims to provide researchers and practitioners with initial evidence on how embedded visualizations impact human-AI reliance in spatial decision-making.
To this end, we present an empirical study (N=32) comparing an AR see-through visualization (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Left) against a 2D minimap (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Right) in a time-critical, AI-assisted spatial decision task. Following previous work on AR visual systems for indoor navigation [59, 91, 90], we selected the AR see-through visualization because it naturally embeds information within a large-scale indoor environment. In the tasks, participants needed to select one of four coffee machines distributed across a two-floor building based on multiple criteria – walking distance and queue length – to obtain coffee as quickly as possible. An AI with 75% accuracy was provided to suggest the optimal target, and both visualizations were experienced using an Apple Vision Pro (AVP), a state-of-the-art AR headset.
Contrary to initial expectations, participants exhibited greater inappropriate reliance on AI with the AR see-through visualization, suggesting that new perceptual and cognitive challenges arise when digital cues are spatially embedded in large or complex physical environments. Further analysis of quantitative data (accuracy, response times, reliance metrics, and self-reported confidence) reveals that the inappropriate reliance primarily manifested as over-reliance – blindly accepting AI suggestions. Through qualitative interviews, we identify several key factors that contributed to the over-reliance, such as perceptual challenges in AR, visual proximity illusions, and heightened trust in “embodied” AI suggestions. On the other hand, our findings also highlight embedded visualizations’ strength in multiple spatial reasoning tasks, including improved spatial mapping (as evidenced by quantitative data) and support for egocentric spatial imagery (as reported by participants). We conclude by outlining the lessons learned, design implications, and promising future directions for better harnessing these AR strengths to foster human–AI collaboration in spatial decision-making.
1 Related Work
Human-AI Decision Making. Human-AI decision making refers to collaborative decision processes where a human decision-maker interacts with an AI system to make a choice [3] . A common goal in these systems is to achieve complementary performance, where the human-AI team performs better than either the human or the AI alone [3] . However, empirical studies have repeatedly found that such complementary performance is rarely observed (e.g., [5, 11]) . One widely reported barrier is inappropriate reliance — the inability of human users to correctly accept AI suggestions when they are right, or reject them when they are wrong [3] .
Unlike trust in AI, which reflects users’ subjective feedback, reliance objectively measures whether a participant accepts or rejects the AI’s suggestion [3, 49] , and has therefore been widely studied. To foster appropriate reliance, researchers have proposed providing explanations of the AI’s suggestions [67]. Unfortunately, several studies found that explanations have often failed to deliver their intended benefits, but actually increase over-reliance – a phenomenon where users uncritically accept AI suggestions, even when they are suboptimal [3]. Drawn on the dual-process model of cognition [12, 22], initial efforts attributed this over-reliance to a lack of cognitive engagement with the explanation, because people often rely on heuristics (Type 1) to judge the AI’s suggestions, rather than engaging in deeper analysis (Type 2). Buçinca et al. [10] introduced cognitive forcing functions that encourage users to think more critically about the AI’s suggestions. Vasconcelos et al. [79] proposed a cost-benefit framework that formalizes when users choose to engage with AI explanations. Recently, Guo et al. [28] developed a decision-theoretic framework to measure AI reliance by separating reliance behavior from cognitive limitations in signal interpretation.
In the context of visualization research, automated methods to assist humans in making decisions with data have long been practiced. While a large body of work has focused on developing visual analytics systems [65], relatively fewer studies have investigated how visualizations affect human-AI collaboration. Among these studies, Yang et al. [93] explored how visual explanations can foster appropriate trust in AI systems; Morrison et al. [51] investigated how different human explanation strategies might impact users’ reliance on AI in visual decision-making; Gaba et al.[24], Wang et al.[83] , and Wall et al. [81] each focused on how interface design influences users’ perceptions of model bias and performance, highlighting the powerful role of visual presentation in shaping trust on AI systems. Recently, Ha et al. [29] found that in AI-guided visual analytics systems, users were more inclined to accept suggestions when completing a more difficult task despite the AI’s lower suggestion accuracy. Zhao et al. [97] and Reyes et al. [63] both investigated how visualizing uncertainty in model outputs influences reliance and trust. Taken together, these studies underscore that visual presentation profoundly influences how users evaluate AI suggestions, rely on them, and ultimately make decisions.
Unlike all these existing works, which evaluate desktop-based human-AI collaboration, our study contributes an initial empirical investigation of the impact of AR visualizations on human-AI collaboration in spatial decision-making tasks. The differences extend beyond display technique (2D monitor vs. AR) to the task environment (desktop vs. physical environment) and to cognitive processes (e.g., general vs. spatial cognition), opening new questions that merit dedicated study.
Visualization-Aided Decision Making. Within the field of visualization, research on decision-making typically falls into two major categories: understanding how humans make decisions when using visualizations (e.g., [54, 15, 4]), and designing interactive systems that support decision-making (e.g., [53]).
The former body of work has drawn from theories in cognitive psychology to model and explain the mechanisms by which visualizations influence decision-making. For example, Padilla et al. [54] proposed a cognitive framework grounded in dual-process theories, illustrating how visualizations engage both intuitive (Type 1) and analytical (Type 2) reasoning processes in different contexts. Bancilhon et al. [2] expanded on this perspective by advocating for the integration of cognitive models in evaluating visualization effectiveness. A parallel line of work investigates how visualizations interact with human cognitive biases. Dimara et al. [15] introduced a task-based taxonomy of cognitive biases relevant to visualization, while Wall et al. [82] explored how interaction traces might mitigate such biases. Bearfield et al. [4] further demonstrated that the same data visualization can elicit diverging interpretations. Research on uncertainty representations also plays an important role in understanding decision-making: Kale et al. [36] showed that design choices, such as emphasizing means, can bias effect size judgments, while Fernandes et al. [23] found that quantile dotplots can significantly improve real-world decisions in uncertain scenarios.
The latter body of work focuses on developing interactive visualization tools to support decision-making. Some of these systems are domain-agnostic and designed to facilitate specific decision processes, such as inspecting rankings [80] and comparing alternatives based on multiple, often conflicting attributes [56] in Multi-Criteria Decision-Making. Another line of work targets domain-specific decision contexts, such as medical treatment planning [39] , urban policy [84] , and more [13] . These systems typically incorporate domain-specific constraints to guide and contextualize the decision process. A recent survey by Oral et al. [53] offers a comprehensive analysis of decision-focused visualization tools and highlights the opportunities and challenges in supporting all stages of the decision-making process.
While decision-making is widely recognized as a core goal of visualization, Dimara and Stasko [16] argue that it remains underrepresented in visualization research and task taxonomies. Brumar et al.[9] recently proposed a typology to better describe decision-making tasks in visualization. Despite this growing interest, decision support for spatial tasks remains an underexplored yet promising area [26]. Our study takes a first step toward investigating spatial decision support, aiming to provide a humble baseline and reference point for future research.
Spatial Cognition and Spatial Decision-Making. Spatial cognition encompasses the mental processes involved in perceiving, encoding, and reasoning about spatial environments [77]. A fundamental aspect of spatial cognition is the use of reference frames—mental models that help individuals locate and orient themselves in space [77, 76] . Two commonly described reference frames are egocentric, where objects are located relative to the observer’s current position and orientation (i.e., “to my left or right”), and allocentric, where the environment is represented independently of the observer’s position (e.g., map-like or object-to-object relationships) [38] . People switch flexibly between these frames, but doing so can introduce significant cognitive load [76] , particularly when tasks require precise mental transformations, such as rotating or scaling spatial representations in mind.
Spatial decision-making requires integrating information about physical environments with external data. To make such decisions with AI support, decision-makers must reconcile digital information with real-world contexts. With traditional 2D map-based visualizations, this process often involves switching between reference frames, which is cognitively demanding and can affect interpreting or verifying AI suggestions. By contrast, AR can display information directly in the physical environment, thereby eliminating the spatial mapping overhead [6]. Yet, to the best of our knowledge, few studies have investigated the effect of AR visualizations on human-AI decision making. Moreover, how to best design effective AR visualizations for spatial decision-making remains an open question. These gaps motivate our empirical investigation into how AR visualizations might influence human-AI collaboration in spatial decision-making.
Situated, Embedded, and AR Visualizations. Different from Virtual Reality (VR), AR can directly visualize information within physical environments, benefiting a broad range of applications that require data-driven or computational support in situ [86, 8]. The visualization community has extensively explored various aspects of AR visualization, including design patterns [69, 42], techniques [98], empirical evaluations [58], and applications [46]. More broadly, the concept of visualizing information in its physical context is known as Situated Visualization [69, 42], for which AR is one realization. Depending on the spatial relationship between the visualization and its physical referent, Willett et al. [87] distinguish between situated (co-located in physical space) and embedded (overlaid or integrated with the referent). Based on this definition, our study compares two AR visualization approaches—a situated 2D minimap and an embedded AR see-through overlay—to investigate whether the embedded approach can foster greater human reliance on AI by eliminating the need to mentally map data from a separate panel onto the physical environment.
AR Visualization for Spatial Decision-Making. Our particular interest lies in leveraging AR to support indoor spatial decision-making under time pressure, a crucial area with high-stakes applications such as emergency evacuation [89, 68], first response operations [96], and security management [73]. Despite its importance, research specifically addressing time-constraint spatial decision-making in AR remains limited. Perhaps the closest related work involves AR-based evacuation assistance, which provides users with dynamic, context-sensitive navigation cues to safely guide them toward exits during emergencies [96]. For instance, Wächter et al. [89] investigated AR-guided evacuation in buildings, showing that adapting AR visualizations to environmental hazards can improve evacuation efficiency and safety. Similarly, Sharma et al. [68] developed AR modules to enhance situational awareness and cognitive mapping during evacuations in buildings.
Beyond emergency scenarios, a broadly related line of work explores AR-based indoor navigation, where users typically follow predefined routes rather than making urgent, real-time spatial decisions. Most AR navigation techniques fall into two categories: annotation-based methods (e.g., arrows) and see-through visualizations. Annotations are arguably the most common AR navigation technique, as seen in applications like Google Maps. Numerous studies (e.g., [43, 17]) have compared annotation-based AR navigation to traditional 2D minimaps across various contexts. On the other hand, AR see-through methods offer benefits like a global view but have received less attention, partly due to technical challenges requiring the development of digital twins of real-world environments. Representative work in this direction includes Xu et al. [91, 90] who conducted comparative studies of AR see-through visualizations for indoor wayfinding, demonstrating that egocentric perspectives significantly enhanced navigation efficiency, reduced cognitive load, and improved spatial awareness.
Drawing inspiration from these works, we adopt AR see-through visualizations to present spatial information within indoor environments. Unlike previous studies, however, our research specifically investigates how AR see-through techniques can support users in making spatial decisions with AI assistance – a topic that has not yet been explored in depth. Our study thus aims to provide an initial reference point for future research into AI-aided decision-making in AR contexts.
2 Study Rationale
2.1 Task Motivation and Abstraction
Spatial decision-making has long been studied under the broader umbrella of wayfinding in spatial cognition (e.g., [21, 85]). Foundational work by Siegel and White [70] conceptualizes wayfinding as a decision process involving three forms of spatial knowledge: landmark, route, and survey. Motivated by real-world applications, our study particularly focuses on target selection (i.e., landmark). In high-stakes scenarios, such as emergency response, indoor search-and-rescue, and security surveillance, spatial decision-making often involves identifying or selecting a target before proceeding with an action (e.g., navigation or intervention). Although AR systems have frequently been investigated for supporting the action phase (e.g., route guidance [90]), their potential to enhance the preceding decision-making step remains underexplored.
  • Space - Large-scale Indoor Environment: These tasks often take place in physical spaces larger than the immediate space around a person (e.g., buildings or open areas) [25, 27]. Tversky [76] characterize such spaces as space of navigation. In this study, we focus on indoor building environments, given their prevalence across numerous application domains.
  • Time - Sensitive: Many high-stakes spatial tasks impose stringent time constraints, requiring users to make optimal decision as soon as possible [52]. In this study, we focus on time pressure scenarios which particularly benefit from decision support systems.
  • Data - Static and Dynamic: Users typically can assess both static information (e.g., building layouts) and dynamic updates (e.g., occupant counts, hazard alerts) from sensors to evaluate each target [1].
  • Decision - Multi-Criteria: These tasks frequently require decisions among multiple targets or locations distributed across rooms, floors, or areas [40]. Success often involves balancing or prioritizing several factors, such as response time, travel distance, and urgency [57].
2.2 Visualization Methods: Minimap vs. AR See-Through
Given that spatial decision-making inherently involves interpreting spatial relationships, an effective decision support tool must present relevant data within its spatial context. Thus, we select map-based visualizations as the foundational interface for our study. We are particularly interested in how different visualization paradigms – traditional 2D vs. embedded – influence users’ decision-making processes. Therefore, we choose two representative visualization methods: a conventional 2D minimap and an AR see-through visualization. Below, we discuss the rationale behind choosing these two methods and briefly highlight alternative visualization approaches.
2D Minimap. 2D minimap is perhaps the most widely adopted map-based visualizations on traditional flat-screens. It presents the spatial environment as a simplified, abstracted, top-down representation, upon which additional data can be annotated (e.g., Google Maps [^1] https://maps.google.com). For indoor environments spanning multiple floors, 2D minimap often employs separate layers for each floor, indicating spatial relationships through key landmarks such as stairways or elevators. 2D minimap is well-established and require minimal training, making them a suitable baseline for comparison. In our study, we implement a minimap (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Left) that shows the spatial distribution of multiple targets and their associated data (i.e., number of people in line) using glyphs.
AR See-Through. Given our task focuses on large-scale indoor environments, we select AR See-Through (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Right) as our embedded visualizations, which overlays a 1:1 scale 3D map on the real-world. It allows users see the interior layouts or objects behind walls, as if they had “X-ray” vision [47], and has been widely used in applications like navigation [91] and construction [72]. For our study, we follow previous similar research [59, 90, 91] that visualize a building’s indoor layout, multiple targets (represented as cubes), and their associated data (i.e., the number of people in line) using virtual avatars.
Other alternatives. We have also considered other map-based alternatives. For example, the 3D minimap provides a three-dimensional representation of the environment and can be implemented on either flat-screen or AR displays. However, 3D minimaps typically rely heavily on interactions such as panning, tilting, zooming, or rotation to fully leverage their advantages. Because our current investigation prioritizes evaluating the impact of visual representations themselves (without the influence of interactive complexity), we intentionally avoid these interactions in our study. Thus, we reserve exploration of interactive 3D minimaps and other advanced visualization methods for future research.
2.3 Visual Decision Supports
A recent survey by Oral et al. [53] highlights a broad spectrum of visual decision support tools, ranging from basic information visualizations to advanced AI-based suggestions [41]. For this study, we specifically choose to implement two fundamental yet representative forms of decision support across both the 2D minimap and AR see-through:
  • Target Visualizations that presents candidate targets and associated real-time data (e.g., occupancy counts, hazard alerts) on the map, without path suggestions or explanatory text.
  • AI-Suggested Optimal Target that highlights a single optimal target based on predefined criteria, with no accompanying explanation. Similar to all AI system, suggestions may not always be accurate. We detail our adjustment of its performance based on prior works [74] and pilot studies in Sec. 3.7.
Following practices established in prior research [79], we deliberately constrain our decision support to these fundamental features for several reasons: First, target and real-time data visualization represent core functionalities broadly used in existing decision-support tools. Second, simple AI suggestions enable direct investigation of user reliance on AI assistance under different visualizations, without the additional cognitive complexity and potential confounding effects introduced by detailed explanations [35].
While more advanced support – such as comparative visualizations, AI-generated explanations, or uncertainty displays – may offer added benefits, we intentionally focus on these fundamental forms to isolate the role of embedded visualizations in spatial decision-making and AI reliance. This controlled setup provides a baseline for future studies exploring richer decision-support designs.
2.4 Hypothesis

### v2

# Can AR Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision-Making? A Comparative Study of AR X-Ray vs. 2D Minimap

@@ 4.2. Spatial Arrangement
  • Floor Relationship – same or cross: All four targets were located on the same floor, which was either the same as the participant’s floor (same) or a different one (cross). We anticipated that this factor would substantially impact the difficulty of estimating walking distances, particularly in the Minimap condition.
  • Proximity – close or far: Targets were categorized as either close or far in direct proximity (instead of walking distance) to the participant’s location. We anticipated that this factor would have a greater impact on task difficulty in the X-ray condition, as distant targets would subtend a smaller retinal angle than closer targets and thus be harder to perceive.
The combination of these two factors resulted in four difficulty levels, ranging from Same+Close (easiest) to Cross+Far (most difficult). Within each trial, to ensure that the appearance of each target was perceptually similar to the user, we used the same difficulty level setting to define each target’s location and placed them all at equal line-of-sight distances from the participant’s location. The participant’s location was then positioned at the center of a circle passing through all four targets. Figure 2 illustrates an example of the four difficulty levels.
Figure 3. The 8 participant locations in the study, distributed across two floors of the building. Participants were instructed to complete 4 trials at each location, following the order indicated by the numbered labels and route. The letter “L” is an abbreviation for location.
[…]
Grouping Arrangements by Participant Location. Ideally, both participant and target locations would vary each trial. However, moving participants between locations would introduce substantial overhead and fatigue. To address this, we grouped the target arrangements by participant location. Specifically, we selected eight participant locations evenly distributed in the building (Figure 3), and for each location, we designed four spatial arrangements of targets, each representing one difficulty level. This resulted in $4\times 8=32$ unique spatial arrangements. An example of one group is shown in Figure 2.
Target Optimality. Target optimality was determined by two metrics: walking distance from the participant and queue length. For simplification, we estimate the total time to obtain coffee from a target using this formula:
$T_{\text{coffee}}(d,n)=\frac{d}{\mathrm{SPEED}}+n\times\mathrm{SERVICE\_TIME}$
where $d$ denotes the walking distance to the target, and $n$ means the number of people in line. $\mathrm{SPEED}$ and $\mathrm{SERVICE\_TIME}$ are constants denote the average walking speed and the average waiting time per person in line, respectively. Based on our pilot studies, we empirically set $\mathrm{SPEED}$ at 1m/s and $\mathrm{SERVICE\_TIME}$ as 15s. While individuals naturally differ in walking speed, participants in our study did not need to physically walk; $\mathrm{SPEED}$ was used for estimation and is thus not expected to introduce substantial bias. These parameters were explicitly communicated to participants in the study.
Since the values of $d$ were fixed by the spatial arrangement of the targets, we manipulated the queue lengths $n$ to differentiate target optimality. Following prior research (Swaroop2024AccuracyTime), each trial was designed to include intentionally varied options: one optimal, two suboptimal, and one worst choice, differing by at least 10 seconds in $T_{\text{coffee}}$. This setup allowed us to evaluate potential participant over-reliance on AI suggestions. Detailed calculations are provided in the supplementary materials.
4.3. Implementation and Apparatus
Following similar work (2021ToTrustOrToThink), a simulated AI was used to provide decision suggestions in applicable trials. Details of the simulated AI setup are provided in Sec. 4.7. We developed the minimap based on high-precision CAD models provided by the building manager. The AR visualizations was developed using ARKit  [^3] https://developer.apple.com/augmented-reality/arkit/ and AR Mesh Manager  [^4] https://docs.unity3d.com/Packages/com.unity.xr.arkit@6.2/manual/arkit-meshing.html to scan the building and construct a 3D digital twin in Unity 6000.0.29f1, with ARAnchor Manager  [^5] https://docs.unity3d.com/Packages/com.unity.xr.arfoundation@6.2/manual/features/anchors/aranchormanager.html ensuring proper registration of the digital model with the physical environment. The system is available at our public repository http://public-when-ready. All trials were completed using an Apple Vision Pro to eliminate device-related differences and provided a high-fidelity immersive experience.
4.4. Participants
We recruited 32 participants (16 male, 16 female; ages 18–28) from our university community via email, word of mouth, and posted flyers. For base compensation, participants received a \$15 Amazon gift card. To incentivize accurate performance, participants earned an additional \$0.20 for each trial in which they selected the optimal target, with the bonus capped at \$20.
To use Apple Vision Pro, participants were required to have normal vision, use contact lenses, or use the prescription inserts provided by us. To minimize potential bias arising from participants’ familiarity with the experimental setting, we exclusively recruited individuals whose prior exposure to the building used in the study was limited. Specifically, 19 participants (9 male, 10 female) had never visited the building, 12 participants (6 male, 6 female) had been in the building between one and five times, and 1 participant (male) had visited the building between five and ten times.
Figure 4. A stimulus workflow from an actual trial. Each trial began with a 20-second countdown. Participants viewed either the X-ray or Minimap visualization while searching for the optimal target. After identifying their choice, participants verbally announced the selected target and rated their confidence on a 1–7 scale. Finally, they physically pointed to the perceived location of the chosen target. For illustration, the figure shows only one trial at location L, but participants completed four trials at each location.
A stimulus workflow from an actual trial. Each trial began with a 20-second countdown. Participants viewed either the X-ray or Minimap visualization while searching for the optimal target. After identifying their choice, participants verbally announced the selected target and rated their confidence on a 1–7 scale. Finally, they physically pointed to the perceived location of the chosen target. For illustration, the figure shows only one trial at location L, but participants completed four trials at each location.
4.5. Procedures
With a within-subjects design, each participant completed trials under four conditions: Minimap+NoAI, Minimap+AI, X-ray+NoAI, and X-ray+AI. The order of conditions was counterbalanced across participants using a Latin square design to minimize order effects and learning biases. Our primary focus was on participants’ behavior in the AI conditions, while the NoAI conditions served as baselines to better understand user decision-making patterns.
32 Participants
$\times$ 4 Conditions: { Minimap+NoAI, Minimap+AI, X-ray+NoAI, X-ray+AI }
$\times$ 2 Participant Location
$\times$ 4 Spatial Arrangements: { Same+Close, Same+Far, Cross+Close, Cross+Far }
1024 total trials (32 per participant)
Figure 5. One example of what a participant experienced during the experiment. Between participants, the order of the 4 conditions (Minimap+NoAI, Minimap+AI, X-ray+NoAI, X-ray+AI) was varied and counterbalanced using a Latin square design. Factors that remained fixed across all participants included the sequence of standing locations (L1-L8) and the order of the four spatial arrangements (Same+Close, Same+Far, Cross+Close, Cross+Far) at each standing location. Within each AI condition, the presentation order of the AI suggestion optimality levels (opt, sub, worst) was randomized, while ensuring a balanced distribution across participants. For illustration, only two examples of AI suggestion presentation orders are shown.
One example of what a participant experienced during the experiment. Between participants, the order of the conditions (rows) was varied and counterbalanced using a Latin square design across 4 conditions (Minimap+AI, Minimap+NoAI, X-ray+AI, X-ray+NoAI). Within each AI condition, the presentation order of the AI suggestion optimality levels (opt, sub, worst) was randomized. Factors that remained fixed across all participants included the sequence of standing locations (L1-L8) and the order of the four spatial arrangements (Same+Close, Same+Far, Cross+Close, Cross+Far) at each standing location.
  • Introduction (10mins): Participants first provided informed consent and were explicitly informed about the performance-based incentives. Next, participants received a detailed tutorial that explained the AVP’s calibration, its usage, and task procedures.
  • Training Tasks (10mins): Participants completed four training trials, one per condition, to familiarize themselves with the task and visualizations. The participant locations and target locations were different in the training trials than in the experimental trials (see supplementary materials). In these trials, after participants made their selections, the optimal target was explicitly highlighted to illustrate optimal decision-making. Participants were encouraged to ask questions and were informed they could withdraw at any point.
  • Actual Tasks (30mins): Participants completed 32 trials, grouped into four blocks, with each block distributed over 2 consecutive locations (see the example in Figure 5). Within each block, one condition was tested (i.e., eight trials per condition). Within each location, they remained physically stationary and completed four trials (one trial per difficulty level). Breaks were allowed during the study.
  • Post-Study Interview (10mins): Participants were interviewed with a list of open-ended questions (attached in supplementary materials).
Detailed Trial Procedure. As illustrated in Figure 4, each trial proceeded through the following steps: The experimenter escorted the participant to the designated starting point and helped them put on the headset. When ready, the experimenter used a controller to initiate the trial, activating the visualization and a 20-second countdown timer. Participants verbally announced their selection, for example, “Target A”, as soon as they identified the optimal target. Then, the experimenter stopped the timer, automatically clearing the visualization. Participants verbally rated their confidence in their choice using a 1 (Strongly unconfident) to 7 (Strongly confident) scale. Lastly, participants physically pointed toward the actual spatial location of their chosen target and looked at their pointing direction. This process was repeated for each trial. We recorded the whole session through audio and first-person view video with participants’ informed consent.
4.6. Measures
For every trial, we documented and calculated the following measures based on prior research (Swaroop2024AccuracyTime; Salimzadeh2024dealWithUncertainty):
  • Decision Accuracy: We defined decision accuracy as the average point score across trials in each condition, ranging from 0 to 1. Following previous research (Swaroop2024AccuracyTime), participants earned points according to their choice quality: optimal choice (1 point), suboptimal choices (0.5 points), and the worst choice (0 points). Each trial contained exactly one optimal target, two suboptimal targets, and one worst target.
  • AI Reliance: Following previous studies (yang2020visual), we categorized participant reliance on AI as either appropriate or inappropriate (Figure 6). We classified reliance as appropriate when participants either followed correct AI suggestions or overrode incorrect ones with better choices. Inappropriate reliance included: (1) over-reliance: accepting suboptimal or worst suggestions; and (2) under-reliance: rejecting correct suggestions in favor of a worse option.
  • Confidence: Participants rated their confidence in their decision on a 7-point scale from 1 (Strongly unconfident) to 7 (Strongly confident).
  • Point Error: Similar to prior research (hu2021comparative; warden2022visual), we assessed whether participants correctly pointed to the spatial location of their selected target based on first-person video recordings. Two co-authors independently coded each response as 0 (correct) or 1 (incorrect) to ensure inter-rater reliability. Inter-rater reliability was evaluated via Cohen’s Kappa ($\kappa=0.898$), suggesting almost perfect agreement according to Landis and Koch’s interpretation (landis1977measurement). Discrepancies were discussed and resolved through consensus. This measure indicates if participants actually knew the spatial location of their selected target.
Figure 6. Classification of reliance as appropriate, over-, and under-reliance based on the alignment between AI suggestions and human decisions. Appropriate reliance on AI was defined as participants accepting optimal suggestions or overriding suboptimal and worst suggestions. Inappropriate reliance included: (1) over-reliance: accepting suboptimal or worst suggestions; and (2) under-reliance: rejecting AI suggestions in favor of a worse option.
Classification of reliance as appropriate, over-, and under-reliance based on the alignment between AI suggestions and human decisions. Appropriate reliance on AI was defined as participants accepting optimal suggestions or overriding suboptimal and worst suggestions. Inappropriate reliance included: (1) over-reliance: accepting suboptimal or worst suggestions; and (2) under-reliance: rejecting AI suggestions in favor of a worse option.
4.7. Using Simulated AI
Similar to prior works (Swaroop2024AccuracyTime; 2021ToTrustOrToThink), we used a simulated AI with an average accuracy of 0.75 to ensure controlled experimentation. In each AI condition, the AI suggested the optimal target in five trials, a suboptimal target in two trials, and the worst target in one trial (Figure 5). This setup yields an average accuracy of 0.75 = $\frac{5\times 1+2\times 0.5+1\times 0}{8}$. The order of these suggestions was randomized and counterbalanced across participants, ensuring that optimal, suboptimal, and worst suggestions were evenly distributed across all trials and participants.
Mitigating Effects of Trust in AI. Following previous work (2021BansalDoesTheWholeExceedItsParts; 2021ToTrustOrToThink), we took steps to neutralize trust effects so that observed differences in reliance could be attributed primarily to the visualizations. Specifically, we avoided anthropomorphizing the AI (Kevin2015TrustInAuto), consistently referring to it as “the AI”. We also described AI outputs as “suggestions” to counteract perfect automation schemas (Mary2002PerceivedUtility) and emphasize that these suggestions could be incorrect. Furthermore, we do not provide decision feedback for each trial (Dietvorst2014AlgorithmAP), or any information about the AI’s accuracy (Yin2019EffectofAccOnTrust).
4.8. Hypothesis